The everyshi package^{∗}

Martin Schröder

Crüsemannallee 3 28213 Bremen Germany martin@oneiros.de

                               PGP-Key:  2048 bit / KeyID 292814E5
                            7E86 6EC8 97FA 2995  82C3 FEA5 2719 090E

2001/05/15

Abstract

This packages provides hooks into \sshipout called \EveryShipout and

     \AtNextShipout analogous to \AtBeginDocument.

Contents

1 Introduction 1

2 Usage 1

3 Options 2 4 Required packages 2

5 Acknowledgements 2

1 Introduction

This package provides the hooks \EveryShipout and \AtNextShipout whose arguments are executed after the output routine has constructed \box255, and before \shipout is called.

An example application for this package would be a package for adding text to the bottom of each page. Such a package does exist: prelim2e[2].

Another application is a package for adding pictures to every page: eso-pic[1].

2 Usage

\EveryShipout{⟨code⟩} declares ⟨code⟩ that is saved internally and executed before each \shipout.

Warning: The ⟨code⟩ is saved globally; there is currently no way to remove it.

\AtNextShipout \AtNextShipout{⟨code⟩} declares ⟨code⟩ that is saved internally and executed

New feature 2001/05/15

just before only the next \shipout.

∗The version umber of this file is v3.00, last revised 2001/05/15.

The name everyshi is a tribute to the 8+3 file-naming convention of certain “operating systems”; strictly speaking it should be everyshipout.

The ⟨code⟩ is executed after \box255 has been constructed by the output routine and can change \box255. \shipout is called after ⟨code⟩.

Repeated use of the commands is permitted: the code in the argument is stored (and executed) in the order of their declarations.

The argument of \AtNextShipout is executed after the argument of \EveryShipout.

3 Options

The package has no options.

4 Required packages

The package does not require any further packages.

5 Acknowledgements

Version 2.0 of everyshi borrows heavily from quire.tex of the Midnight Macros by Marcel R. van der Goot (marcel@cs.caltech.edu). The pointer to quire was provided by Lothar Meyer-Lerbs (TEXSatz@zfn.uni-bremen.de). As usual Rebecca Stiels improved the quality of this documentation.

References

[1] Rolf Niepraschk. The eso-pic package. CTAN:tex-archive/macros/latex/

   contrib/supported/ms/contrib/eso-pic.dtx. L^{A}TEX 2ε package.

[2] Martin Schröder. The prelim2e package.

   CTAN:tex-archive/macros/latex/contrib/supported/ms/prelim2e.dtx.

L^{A}TEX 2ε package.